Ferris Company began January with 9,000 units of its principal product. The cost of each unit is $8. Merchandise transactions for the month of January are as follows:

Purchases
Date of Purchase Units Unit Cost* Total Cost
Jan. 10 6,000 $ 9 $ 54,000
Jan. 18 9,000 10 90,000
Totals 15,000 144,000


* Includes purchase price and cost of freight.

Sales
Date of Sale Units
Jan. 5 5,000
Jan. 12 3,000
Jan. 20 6,000
Total 14,000


10,000 units were on hand at the end of the month.

Required:
1. Calculate January's ending inventory and cost of goods sold for the month using FIFO, periodic system.

2. Calculate January's ending inventory and cost of goods sold for the month using LIFO, periodic system.

3. Calculate January's ending inventory and cost of goods sold for the month using FIFO, perpetual system.

4. Calculate January's ending inventory and cost of goods sold for the month using Average cost, periodic system.

5. Calculate January's ending inventory and cost of goods sold for the month using Average cost, perpetual system. (Round average cost per unit to 4 decimal places. Enter sales with a negative sign.)
Answer #1

Pictures have been uploaded for first three parts. Please refer to a brief description of the answers below:

1. FIFO calculates inventory at the price of the latest lot of goods.

Closing inventory = 9000*10 + 1000*9 = $ 99,000

COGS = Opening inventory + Purchases - Closing Inventory = 72000 + 144000 - 99000 = $ 117,000

2. LIFO calculates inventory at the price of the oldest lot.

Closing Inventory = 9000*8 + 1000*9 = $ 81,000

COGS = 72000 + 144000 - 81000 = $ 135,000

3. The 3rd part is solved in an easy table format. In case it is not understandable, please leave a comment to make me aware of it.

4. Units available for sale = Opening Inventory + Purchases = 9000+15000 = 24,000 units

Value of goods available for sale = 72,000+144,000 = $ 216,000

Average cost = Total value / units = 216,000/24,000 = $ 9

Closing Inventory = 10,000*9 = $ 90,000

COGS = 72000+144000-90000 = $ 126,000

5. Closing Inventory = 10,000*9.3875 = $ 93,875

COGS = $ 122,125

Average caet method SPAGE balance Date Rececipte Isculs Unite Rati amt Units Rate tunits Rate 9000 Jan 1 Tam 5 Jom 10 5000 3.

feriodic FIPO 10,000 unite laft will be IPO method picid at $1o and to Inwentoy which m ths wits aie ustodt the lot peeds einSPP PERPETUAL SYSTEM DATE PAGE FIFO method Date Reststs Balance Tan Jan 5 9000 7 000 5000 4ou0 39000 G000 SYOO0 4000 32000 60